# On Speeding Up Language Model Evaluation

### v2

@@ Abstract
Developing prompt-based methods with Large Language Models (LLMs) requires making numerous decisions, which give rise to a combinatorial search problem. For example, selecting the right pre-trained LLM, prompt, and hyperparameters to attain the best performance for a task typically necessitates evaluating an expoential number of candidates on large validation sets. This exhaustive evaluation can be time-consuming and costly, as both inference and evaluation of LLM-based approaches are resource-intensive. Worse, a lot of computation is wasted: Many hyper-parameter settings are non-competitive, and many samples from the validation set are highly correlated - providing little or no new information. So, if the goal is to identify the best method, it can be done far more efficiently if the validation samples and methods are selected adaptively. In this paper, we propose a novel method to address this challenge. We lean on low-rank matrix factorization to fill in missing evaluations and on multi-armed bandits to sequentially identify the next (method, validation sample)-pair to evaluate. We carefully assess the efficacy of our approach on several competitive benchmark problems and show that it can identify the top-performing method using only 5-15% of the typically needed resources—resulting in a staggering 85-95% LLM cost savings.
@@ 1 Introduction
Figure 1: Comparison of cost for finding the best model or prompt between our proposed algorithms (UCB-E, UCB-E-LRF) and Full Evaluation on two datasets. Left: AlpacaEval [44] data as of May 20, 2024. The data includes 153 models evaluated on 805 instructions. The response on each instruction for each model is compared against a baseline response generated by GPT4-turbo, with GPT4-turbo being the LLM judge [66]. The monetary saving is calculated based on the inference cost of proprietary models among the 153 models and GPT4-turbo judging cost. Right: GSM8K [18] prompt engineering data. The data includes 205 zero-shot prompts for Mistral-7B generated by GPT4 evaluated on 784 questions (10%) of GSM8K training set. The time saving is calculated based on inference speed of Mistral-7B using the Huggingface library with one Nvidia A6000 GPU. The overhead computation time for UCB-E and UCB-E-LRF is 2.4 and 142.6 seconds, respectively.
@@
Large language models (LLMs) have demonstrated remarkable proficiency in diverse tasks such as question answering, machine translation and mathematical reasoning [5, 16, 43]. They are employed across numerous applications, ranging from automated customer support to content generation [45]. As the development of new LLMs continues, practitioners find themselves with a plethora of options, selecting the optimal model, prompt [40], or hyperparameters for their specific needs from hundreds available. For instance, initiatives like the Chatbot Arena [15] actively maintain nearly 100 models for benchmarking on user-specified open-ended queries. Similarly, the AlpacaEval [44] project benchmarks over 200 models against a diverse set of 805 questions.
@@
Querying an LLM is resource-intensive, and therefore such extensive evaluations require significant investments in time, compute and financial resources. Figure 1 (left) shows the estimated cost of a Full Evaluation of the 153 models officially included in AlpacaEval [44] as of May 20, 2024 to be almost 800 USD. Similarly, 78 Nvidia A6000 GPU hours are needed for evaluating 205 zero-shot prompts on 784 GSM8K [18] questions using Mistral-7B [32], Figure 1 (right).
Although it is common practice to exhaustively evaluate all validation samples of a given task for each method, practitioners often only care about the final aggregated performance rankings. In other words, the goal is to identify the top-performing methods, discarding the lower-ranked alternatives. Consequently, evaluating on all validation samples is often unnecessarily costly and wasteful.
@@
In this paper, we explore a limited-budget setting. Given a set of methods we aim to identify the one that has the best average performance among a set of validation examples — given a fixed evaluation budget. A simple baseline is to evenly split the budget for each method. However, this can be very inefficient as it would spend a lot of the budget on obvious non-performers and too little effort on teasing out what the best methods are. Intuitively, methods that consistently underperform other methods are unlikely to end up as the best. Therefore, a more efficient budget allocation strategy is to spend less on low-performing methods and more on promising ones. Multi-armed bandit algorithms enable this dynamic allocation by actively selecting the next method-example pair to evaluate based on the results of previous evaluations.
@@
We propose two active selection algorithms UCB-E and UCB-E-LRF. Our first algorithm is an extension of the classical UCB-E [2] to solve the multi-armed bandit problem. It estimates the upper confidence bound (UCB) to guide the selection of the method which is paired with a randomly chosen example for the next evaluation in order to efficiently estimate the best method. This algorithm enjoys a strong theoretical guarantee that the chance of selecting the best arm converges to 100% by an exponential decay of the number of evaluations. Our second algorithm, UCB-E-LRF, leverages the fact that in practice many validation samples are similar, and their results are highly correlated. In other words, if a method performs well on one sample, it will predictably also perform well on similar samples. We can therefore speculatively fill in those results and evaluate the method on a sample that is less correlated. More formally, imagine we have $n$ methods and $m$ data points and store the evaluation results in a $(n\times m)$ matrix $E$. We can leverage the fact that this matrix $E$ is often low-rank, or can be approximated well by a low-rank matrix; evident from Table 3 and Figure 6 in Appendix A. We therefore regard the evaluation process as a matrix completion problem, where we predict the remaining unobserved $E_{i,j}$ entries by low-rank factorization (LRF) and prioritize evaluations of entries with large prediction uncertainties. By deploying this approach in addition to UCB-E, UCE-E-LRF actively selects both the method and example to evaluate, which leads to even bigger budget savings in settings with correlated validation samples.
@@
We evaluate the efficacy of these two algorithms, in addition to common non-active baselines, in a number of practical settings. Our empirical analysis shows that our two active selection algorithms are much better than all baselines. As predicted, the benefits of active selection are larger if the top methods are separated by a bigger performance gap. Notably, we observe that UCB-E works best for such easier settings; and UCB-E-LRF shines in harder settings, when the performance differences between methods are more subtle.
2 Preliminaries
2.1 Evaluation Workflow in Large Language Model Applications
@@
BOLD [22] GLUE [58] HellaSwag [64]
SQuAD [52] TriviaQA [34] WinoGrande [1]
CNN/Dailymail [48] Newsroom [25]
XSUM [49]
Open-ended QA Google NQ [41] HotpotQA [62] QASPER [19]
AlpacaEval [44] Chatbot Arena [15]
MT-Bench [66]
@@
APPS [27] Arc [17] GPQA [53] GSM8K [18]
MATH [29] MMLU [28] PIQA [3]
@@
    Inference. Given a dataset (see Table 1), the output is generated by a specific LLM-based method. Here, methods can differ based on their LLM architecture (architecture search), a prompt (for prompt engineering [60]), or specific decoding parameters (e.g. temperature, decoding strategies), or hyper-parameters.
  2. 2.
@@
    Scoring. The outputs from different methods are scored with a scoring function i.e. metric. The scoring function can either be rule-based (exact string match, BLEU [50], ROUGE [46]), LLM-based (BERTScore [65], LLM judge [66]), or human-based. Depending on the task and dataset format, researchers have employed different types of scoring functions. In Table 1, we classify a selected group of commonly used datasets according to the task type and scoring function applied to them. We group LLM-based and human-based scoring together since they are usually considered alternatives to each other and have been shown to have relatively high correlation [44].
  3. 3.
    Performance aggregation. The performance of each method is aggregated across the dataset, often with a simple average over all examples in the dataset, although many alternative aggregations are possible.
@@
Both, inference and scoring do often rely heavily on LLM-based scoring through black-box APIs [42]—inducing large costs and resource consumption. However, in many practical scenarios, we are only interested in identifying the best method among all methods. For example, for prompt engineering and hyperparameter tuning, knowing which method (prompt / configuration) is the best is usually sufficient for next round of iteration. Identifying exactly which method is ranked 84 out of 100 is irrelevant and frankly, a waste of resources. We therefore claim that evaluating all methods on all examples of a dataset is excessive for this purpose, which motivates the research question how to identify the best method with as few evaluations as necessary.
2.2 Notation and Problem Formulation
Suppose we have a set of examples $\mathcal{X}=\{x_{1},\ldots,x_{m}\}$ and a set of methods =$\mathcal{F}=\{f_{1},\ldots,f_{n}\}$, such that $f_{i}:\mathcal{X}\to\mathcal{Y}$. Let $e:\mathcal{Y}\to[0,1]$ denote our scoring function and without loss of generality, we assume $e(f(x))>e(f^{\prime}(x))$ indicates $f$ has better performance than $f^{\prime}$ on the example $x$. Suppose $E\in[0,1]^{n\times m}$ is the underlying scoring matrix for a given problem $(\mathcal{F},\mathcal{X},e)$ by defining $E_{ij}:=e(f_{i}(x_{j}))$. The score of a given method $f_{i}$ is defined as $\mu_{i}=\frac{1}{m}\sum_{j=1}^{m}E_{ij}$ and we want to find the best method $i^{*}=\arg\max_{i}\mu_{i}.$ Formally, given a finite evaluation budget $T$, we want to study the evaluation algorithm ${\mathcal{A}}(\mathcal{F},\mathcal{X},T)$ and maximize $\mathbb{P}_{\mathcal{A}}(\mathcal{A}(T,\mathcal{F},\mathcal{X})=i^{*})$, the probability of returning the best method $i^{*}$ within budget $T$.
@@ 2.2 Notation and Problem Formulation
We introduce the following additional notation to allow for a clear description of the algorithms. Let $E^{\rm obs}\in\left([0,1]\cup\{?\}\right)^{n\times m}$ denote the partially observed scoring matrix, and $O\in\{0,1\}^{n\times m}$ denote the observation matrix where $O_{ij}$ indicates whether the method-example pair $(f_{i},x_{j})$ has been observed. Finally let us denote our estimate of $E$ by $\hat{E}$.
@@ 3 Algorithms
One simple baseline for designing the evaluation algorithm $\mathcal{A}$ is: uniformly sample and evaluate $T$ examples from $\mathcal{X}$, for each method $f_{i}\!\in\!\mathcal{F}$, estimate the performance $\hat{\mu}_{i}=\frac{1}{\sum_{j=1}^{m}O_{ij}}\sum_{j=1}^{m}O_{ij}\cdot E^{\rm obs}_{ij}$, and pick the method $f_{\hat{i}^{*}}$ with the highest estimated mean $\hat{\mu}_{\hat{i}^{*}}$ as the prediction for the best method. This baseline ignores any information gathered in the process of sequentially evaluating examples, and, as expected, can be very inefficient. We will see in the experiments that for some datasets, this baseline needs to evaluate at least $90\%$ of all method-example pairs to predict the best method correctly with high probability. In the following section, we will introduce two adaptive algorithms that actively select the next method-example pair to evaluate based on previous results; Figure 2 illustrates this idea on a high level for prompt tuning. The methods in the figure are different prompts, and the examples are simple math problems.
@@ 3.1 Algorithm 1 – UCB-E
We address this limitation by adaptive selecting example $t$ based on our previous $t-1$ observations until out budget $T$ is expended. Our first algorithm UCB-E ($\mathcal{A}_{\rm ue}$; Algorithm 1) is a simple extension of the classic multi-arm bandit algorithm UCB-E [2]. For every method $f_{i}$, at every step $t$, we estimate the upper confidence bound,
$B_{i}=\frac{\sum_{j=1}^{m}O_{ij}\cdot E^{\rm obs}_{ij}}{\sum_{j=1}^{m}O_{ij}}+\sqrt{\frac{a}{\sum_{j=1}^{m}O_{i,j}}}.$
In the subsequent iteration we pick the method $f_{i}$ with the largest upper confidence bound (ties are resolved by uniform random sampling), and then sample an example $x_{j}$ uniformly at random from the set of examples not yet observed, that is the set $\{j\in[m]|O_{ij}=0\}$. We run the inference procedure for $f_{i}(x_{j})$ and score the result, obtaining $E^{obs}_{ij}=e(f_{i}(x_{j}))$.
@@
The UCB-E algorithm efficiently addresses the limitations of the simple baseline by minimizing evaluations on suboptimal methods. For a bad method (i.e., when the gap $\mu_{i^{*}}-\mu_{i}$ is large), only a few evaluations (typically proportional to $\frac{1}{(\mu_{i^{*}}-\mu_{i})^{2}}$) are needed to ensure that its upper confidence bound will never exceed $\mu_{i^{*}}$. Consequently, the UCB-E algorithm will not select this method again, thus saving evaluation budget and focusing on more promising methods.
@@
1:  The upper confidence bounds $B:=\{+\infty\}^{n}$, the observation matrix $O:=\{0\}^{n\times m}$, the observed scoring matrix $E^{\rm obs}:=\{?\}^{n\times m}$.
@@
3:     Select: Draw uniformly at random $i\in\arg\max_{k\,\mid\,(\sum_{j=1}^{m}O_{kj})\neq m}B_{k}$; Draw uniformly at random $j\in\{k\in[m]|O_{ik}=0\}$.
4:     Evaluate: Run inference for the method-example pair $(f_{i},x_{j})$, score the result, and receive $e(f_{i}(x_{j}))$; $E^{\rm obs}_{ij}\leftarrow e(f_{i}(x_{j}))$.
5:     Update: $O_{i,j}\leftarrow 1$; $B_{i}\leftarrow\frac{\sum_{j=1}^{m}O_{ij}\cdot E^{\rm obs}_{ij}}{\sum_{j=1}^{m}O_{ij}}+\sqrt{\frac{a}{\sum_{j=1}^{m}O_{ij}}}$.
@@
Return: $\hat{i}^{*}=\arg\max_{i}\frac{\sum_{j=1}^{m}O_{ij}\cdot E^{\rm obs}_{ij}}{\sum_{j=1}^{m}O_{ij}}$
@@
We state the theoretical guarantee of UCB-E in our context, which is a corollary of the theorem for UCB-E in the original multi-arm bandit setting [2].
@@
Define $H_{1}=\sum_{i=1,i\neq i^{*}}^{n}\frac{1}{(\mu_{i}-\mu_{i}^{*})^{2}}$ and suppose $a=\frac{25}{36}\frac{T-n}{H_{1}}$, $\mathbb{P}_{\mathcal{A}_{\rm ue}}(\mathcal{A}_{\rm ue}(T,\mathcal{F},\mathcal{X};a)=i^{*})\geq 1-2Tn\exp\left(-\frac{T-n}{18H_{1}}\right)$.
Comment. $H_{1}$ in the corollary indicates the hardness of the problem specified by $(\mathcal{F},\mathcal{X},e)$. The larger the gaps between the highest and all other scores are, the higher the probability that UCB-E will pick the best method given a fixed budget $T$. In our experiments, we will observe that the datasets with smaller $H_{1}$ tend to have a higher chance of predicting $i^{*}$ correctly, given the same evaluation budget $T$.

### v3

@@ Abstract
Developing prompt-based methods with Large Language Models (LLMs) requires making numerous decisions, which give rise to a combinatorial search problem over hyper-parameters. This exhaustive evaluation can be time-consuming and costly. In this paper, we propose an adaptive approach to explore this space. We are exploiting the fact that often only few samples are needed to identify clearly superior or inferior settings, and that many evaluation tests are highly correlated. We lean on multi-armed bandits to sequentially identify the next (method, validation sample)-pair to evaluate and utilize low-rank matrix factorization to fill in missing evaluations. We carefully assess the efficacy of our approach on several competitive benchmark problems and show that it can identify the top-performing method using only 5-15% of the typical resources—resulting in 85-95% LLM cost savings. Our code is available at https://github.com/kilian-group/banditeval.
@@ 1 Introduction
Figure 1: Cost comparison for finding the best model or prompt between our proposed algorithms (UCB-E, UCB-E-LRF) and Full Evaluation on two datasets. The overhead computation time for UCB-E and UCB-E-LRF is 2.4 and 142.6 seconds, respectively. See text for details.
@@
Large language models (LLMs) have demonstrated remarkable proficiency in diverse tasks such as question answering, machine translation, and mathematical reasoning (Brown et al., 2020; Chowdhery et al., 2023; Lewkowycz et al., 2022). They are employed across numerous applications, ranging from automated customer support to content generation (Liang et al., 2024). As the development of new LLMs continues, practitioners find themselves with a plethora of options, selecting the optimal model, prompt (Kojima et al., 2022), and hyperparameters for their specific needs.
@@
Querying an LLM is resource-intensive, and extensive evaluation requires significant investment in capital, time, and compute. Figure 1 (left) shows the estimated cost of a Full Evaluation of the 153 models officially included in AlpacaEval (Li et al., 2023) as of May 20, 2024 to be almost 800 USD. GPT4-turbo is the LLM judge (Zheng et al., 2024) and the monetary cost is calculated based on the inference cost of only proprietary models and GPT4-turbo judging cost. Similarly, 78 Nvidia A6000 GPU hours are needed for evaluating 205 zero-shot prompts on 784 GSM8K (Cobbe et al., 2021) questions using Mistral-7B (Jiang et al., 2023), Figure 1 (right).
In this paper we investigate the familiar setting where given a set of methods we aim to identify the best performing across a set of validation examples — given a fixed evaluation budget. For example, the different methods could be LLM prompts or hyperparameter settings and the validation examples could be translation tasks or K12 math problems. The budget could be a fixed monetary sum or GPU time. We argue that this setting is extremely common among LLM applications, where the practitioner is primarily interested in identifying the single best method to deploy.
@@
A simple baseline is to evenly split the budget for each method. However, this can be very inefficient as it would spend a lot of the budget on obvious non-performers and too little effort on identifying the very best method. Intuitively, methods that consistently underperform other methods are unlikely to end up as the best. Therefore, a more efficient budget allocation strategy is to spend less on low-performing methods and more on promising ones. Multi-armed bandit algorithms enable this dynamic allocation by actively selecting the next method-example pair to evaluate based on the results of previous evaluations.
@@
We propose two active selection algorithms UCB-E and UCB-E-LRF. Our first algorithm is an extension of the classical UCB-E (Audibert and Bubeck, 2010) to solve the multi-armed bandit problem. It estimates the upper confidence bound (UCB) to guide the selection of the method which is paired with a randomly chosen example for the next evaluation in order to efficiently estimate the best method. This algorithm has a strong theoretical guarantee that the probability of selecting the best arm approaches 100% with an exponential convergence in the number of evaluations. Our second algorithm, UCB-E-LRF (for Low-Rank Factorization), leverages the fact that in practice many validation samples are similar, and their results are highly correlated. In other words, if a method performs well on one sample, it will predictably also perform well on similar samples. We can therefore speculatively fill in those results and evaluate the method on a sample for which the performance is more uncertain. In our method, we capture this insight through a low-rank approximation of the evaluation matrix. By deploying this approach in addition to UCB-E, UCE-E-LRF actively selects both the method and example to evaluate, which leads to even bigger budget savings in settings with correlated validation samples.
@@
We evaluate the efficacy of these two algorithms, in addition to common non-active baselines, in a number of practical settings. Our empirical analysis shows that our two active selection algorithms are much better than all baselines. Not surprisingly, the benefits of active selection are larger if the top methods are separated by a bigger performance gap. Notably, we observe that UCB-E works best for such easier settings; and UCB-E-LRF shines in harder settings, when the performance differences between methods are more subtle.
2 Notation and Problem Formulation
@@
BOLD Dhamala et al. (2021) GLUE Wang et al. (2019)
HellaSwag Zellers et al. (2019) SQuAD Rajpurkar et al. (2016)
TriviaQA Joshi et al. (2017) WinoGrande ai2 (2019)
CNN/Dailymail Nallapati et al. (2016)
Newsroom Grusky et al. (2018)
XSUM Narayan et al. (2018)
Open-ended QA
Google NQ Kwiatkowski et al. (2019) HotpotQA
Yang et al. (2018) QASPER Dasigi et al. (2021)
AlpacaEval Li et al. (2023)
Chatbot Arena Chiang et al. (2024)
MT-Bench Zheng et al. (2024)
@@
APPS Hendrycks et al. (2021a) Arc Clark et al. (2018)
GPQA Rein et al. (2023) GSM8K Cobbe et al. (2021)
MATH Hendrycks et al. (2021c) MMLU Hendrycks et al. (2021b)
PIQA Bisk et al. (2020)
@@
Given a dataset of examples $\mathcal{X}=\{x_{1},\ldots,x_{n}\}$, outputs are generated by specific LLM-based methods =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}$. Here, methods can differ based on their LLM architecture (architecture search), a prompt (prompt engineering (Kojima et al., 2022)), or specific decoding parameters (e.g. temperature, decoding strategies), or other hyper-parameters.
Scoring.
@@
The outputs from different methods are scored with a scoring function $s:\mathcal{Y}\to[0,1]$. Without loss of generality, we assume $s(f(x))>s(f^{\prime}(x))$ indicates $f$ has better performance than $f^{\prime}$ on an example $x$. The scoring function can either be computational (exact string match, BLEU (Papineni et al., 2002), ROUGE (Lin, 2004)), LLM-based (BERTScore (Zhang et al., 2019), LLM judge (Zheng et al., 2024)), or human annotators. Depending on the task and dataset format, researchers have employed different types of scoring functions. In Table 1, we classify a selected group of commonly used datasets according to the task type and scoring function applied to them. We group LLM-based and human-based scoring together since they are usually considered alternatives to each other and have been shown to have relatively high correlation (Li et al., 2023).
Performance aggregation.
Once the underlying scoring matrix $S\in[0,1]^{m\times n}$ is computed as $S_{ij}:=s(f_{i}(x_{j}))$, the aggregate performance of method $f_{i}$ is its average score across all validation samples, $\mu_{i}=\frac{1}{n}\sum_{j=1}^{n}S_{ij}$. Alternative aggregations are possible but not considered in this work.
@@
Both inference and scoring often rely heavily on LLM-based scoring through black-box APIs (Lambert et al., 2024)—inducing large costs and resource consumption. However, in many practical scenarios, we are only interested in identifying the best method among all methods. For example, for prompt engineering and hyperparameter tuning, knowing which method (prompt / configuration) is the best is usually sufficient for the next round of iteration. Identifying exactly which method is ranked 84 out of 100 is irrelevant and a waste of resources. We therefore claim that evaluating all methods on all examples of a dataset is excessive for this purpose, which motivates the research question of how to identify the best method $i^{*}=\arg\max_{i}\mu_{i}$ with limited evaluations.
@@
Formally, given a finite evaluation budget $T$, we want an algorithm ${\mathcal{A}}(T,\mathcal{F},\mathcal{X})$ that maximizes $\mathbb{P}_{\mathcal{A}}(\mathcal{A}(T,\mathcal{F},\mathcal{X})=i^{*})$, the probability of returning the best method $i^{*}$ within budget $T$. We introduce the following additional notation to allow for a clear description of the algorithms. Let $S^{\rm obs}\in\left([0,1]\cup\{?\}\right)^{m\times n}$ denote the partially observed scoring matrix, and $O\in\{0,1\}^{m\times n}$ denote the observation matrix where $O_{ij}$ indicates whether the method-example pair $(f_{i},x_{j})$ has been observed. Finally, let us denote our estimate of $S$ by $\hat{S}$. For ease of reference, We also list the notations used throughout the paper in Appendix E Table 4.
@@ 3 Algorithms
One simple baseline for designing the evaluation algorithm $\mathcal{A}$ is: uniformly sample and evaluate $T$ examples from $\mathcal{X}$, for each method $f_{i}\!\in\!\mathcal{F}$, estimate the performance $\hat{\mu}_{i}=\frac{1}{\sum_{j=1}^{n}O_{ij}}\sum_{j=1}^{n}O_{ij}\cdot S^{\rm obs}_{ij}$, and pick the method $f_{\hat{i}^{*}}$ with the highest estimated mean $\hat{\mu}_{\hat{i}^{*}}$ as the prediction for the best method. This baseline ignores any information gathered in the process of sequentially evaluating examples, and, as expected, can be very inefficient. We will see in the experiments that for some datasets, this baseline needs to evaluate at least $90\%$ of all method-example pairs to predict the best method correctly with high probability. In the following section, we will introduce two adaptive algorithms that actively select the next method-example pair to evaluate based on previous results; Figure 2 illustrates this idea on a high level for prompt tuning. The methods in the figure are different prompts, and the examples are simple math problems.
@@ 3.1 Algorithm 1 – UCB-E
We address this limitation by adaptively selecting example $t$ based on our previous $t-1$ observations until our budget $T$ is expended. Our first algorithm UCB-E ($\mathcal{A}_{\rm ue}$; Algorithm 1) is a simple extension of the classic multi-arm bandit algorithm UCB-E (Audibert and Bubeck, 2010). For every method $f_{i}$, at every step $t$, we estimate the upper confidence bound,
$B_{i}=\frac{\sum_{j=1}^{n}O_{ij}\cdot S^{\rm obs}_{ij}}{\sum_{j=1}^{n}O_{ij}}+\sqrt{\frac{a}{\sum_{j=1}^{n}O_{i,j}}}.$
Here, $a$ is a hyperparameter that controls the tradeoff between exploration and exploitation, where a larger $a$ value encourages more exploration. In the subsequent iteration, we pick the method $f_{i}$ with the largest upper confidence bound (ties are resolved by uniform random sampling), and then sample an example $x_{j}$ uniformly at random from the set of examples not yet observed, that is the set $\{j\in[n]|O_{ij}=0\}$. We run the inference procedure for $f_{i}(x_{j})$ and score the result, obtaining $S^{obs}_{ij}=s(f_{i}(x_{j}))$.
@@
The UCB-E algorithm addresses the limitations of the simple baseline by minimizing evaluations on suboptimal methods. For a bad method (i.e. $\mu_{i^{*}}-\mu_{i}$ is large), only a few evaluations (typically proportional to $\frac{1}{(\mu_{i^{*}}-\mu_{i})^{2}}$) are needed to ensure that its upper confidence bound will never exceed $\mu_{i^{*}}$. Consequently, the UCB-E algorithm will not select this method again, thus saving the evaluation budget and focusing on more promising methods.
@@
1:  The upper confidence bounds $B:=\{+\infty\}^{m}$, the observation matrix $O:=\{0\}^{m\times n}$, the observed scoring matrix $S^{\rm obs}:=\{?\}^{m\times n}$.
@@
3:     Select: Draw $i\in\arg\max_{k\,\mid\,(\sum_{j=1}^{n}O_{kj})\neq n}B_{k}$; Draw uniformly at random $j\in\{k\in[n]|O_{ik}=0\}$.
4:     Evaluate: Run inference for the method-example pair $(f_{i},x_{j})$, score the result, and receive $s(f_{i}(x_{j}))$; $S^{\rm obs}_{ij}\leftarrow s(f_{i}(x_{j}))$.
5:     Update: $O_{ij}\leftarrow 1$; $B_{i}\leftarrow\frac{\sum_{j=1}^{n}O_{ij}\cdot S^{\rm obs}_{ij}}{\sum_{j=1}^{n}O_{ij}}+\sqrt{\frac{a}{\sum_{j=1}^{n}O_{ij}}}$.
@@
Return: $\hat{i}^{*}=\arg\max_{i}\frac{\sum_{j=1}^{n}O_{ij}\cdot S^{\rm obs}_{ij}}{\sum_{j=1}^{n}O_{ij}}$
@@
We state the theoretical guarantee of UCB-E in our context, which is a corollary of the theorem for UCB-E in the original multi-arm bandit setting (Audibert and Bubeck, 2010).
@@
Define $H_{1}=\sum_{i=1,i\neq i^{*}}^{m}\frac{1}{(\mu_{i}-\mu_{i}^{*})^{2}}$ and suppose $a=\frac{25}{36}\frac{T-m}{H_{1}}$, $\mathbb{P}_{\mathcal{A}_{\rm ue}}(\mathcal{A}_{\rm ue}(T,\mathcal{F},\mathcal{X};a)=i^{*})\geq 1-2Tm\exp\left(-\frac{T-m}{18H_{1}}\right)$.
Comment. $H_{1}$ in the corollary indicates the hardness of the problem specified by $(\mathcal{F},\mathcal{X},s)$. Intuitively, the more methods that have a smaller performance gap with the best method, the harder it is to identify which method is the best. We observe that datasets with smaller $H_{1}$ tend to have a higher chance of predicting $i^{*}$ correctly, given the same evaluation budget $T$.